# An Evidence-Based Post-Hoc Adjustment Framework for Anomaly Detection Under Data Contamination

### v1

@@ 4.1 Extension to score-based anomaly detection
Following energy-based model (EBM) formulations (LeCun et al.,, 2006), we can define the associated contaminated PDF as
@@ 4.2 An illustrative example
To illustrate the effect of EPHAD, we use a toy dataset inspired by Qiu et al., (2022). The dataset is generated using a two-dimensional mixture model comprising three Gaussian components: $c_{1}:={\mathcal{N}}(\mu_{1},\Sigma_{1}),c_{2}:={\mathcal{N}}(\mu_{2},\Sigma_{2}),c_{3}:={\mathcal{N}}(\mu_{3},\Sigma_{3})$. Here, each component follows a Gaussian distribution ${\mathcal{N}}(\mu,\Sigma)$ with mean $\mu$ and covariance $\Sigma$. Normal samples are drawn from $f^{+}_{X}=c_{1}$, with $\mu_{1}=[1,1]^{T}$ and $\Sigma_{1}=0.07~\mathbf{I}_{2}$. Anomalous samples are drawn from a mixture distribution $f^{-}_{X}:=0.5c_{2}+0.5c_{3}$ where $\mu_{2}=[-0.25,2.5]^{T},\mu_{3}=[-1,0.5]^{T}$ and $\Sigma_{2}=\Sigma_{3}=0.03~\mathbf{I}_{2}$. The extended implementation details is provided in Appendix B.2. Using this setting, we create a contaminated dataset consisting $100$ data points. We compare the baseline DeepSVDD (Ruff et al.,, 2018) across three configurations as illustrated in Figure 1: (i) “Blind”, (ii) “Refine”, and (iii) with EPHAD. We refer to the baseline model that treats all samples as normal as “Blind”, while “Refine” denotes a model that iteratively filters out suspected anomalies during training. As an evidence function in EPHAD, LOF (Breunig et al.,, 2000) is computed on test samples at test time. The results in Figure 1 demonstrate that the “Blind” configuration mistakenly considers all anomalies as normal. The “Refine” configuration improves performance by filtering out a subset of anomalies. Finally, EPHAD establishes a clearer boundary around normal samples.
@@ 4.3 Determining the temperature parameter $\beta$
As previously discussed, EPHAD has only a single hyperparameter, $\beta$, which controls the trade-off between reliance on the original AD model and the evidence function $T(x)$. A straightforward approach to selecting $\beta$ would involve evaluating the AD performance of the prior and $T(x)$ individually on a validation set and choosing $\beta$ accordingly. However, this strategy introduces additional computational overhead at test time and requires access to a labelled validation set of sufficient size to ensure reliable performance estimation – conditions often impractical in real-world deployments. To address this limitation, we propose an adaptive extension of our approach, termed EPHAD-Ada, that determines the optimal $\beta$ in an unsupervised manner using only test data at test time. This adaptation is inspired by the principle of Entropy Minimisation (EM) (Press et al.,, 2024), a widely-used technique in test-time adaptation (Xiao and Snoek,, 2024). Motivated by the observation from Wang et al., (2021) that models tend to be more accurate when predictions are made with high confidence, we apply it to compute the hyperparameter $\beta$. Specifically, the computation of $\beta$ depends on the entropy of the inlier probabilities derived from the scores of both the original model and the evidence function.
@@
where $p_{s}:={\mathbb{P}}(S\leq s)$. Since $p_{s}$ is unknown in practice, we follow the approach of Perini et al., (2021) and treat it as a random variable $P_{s}$ with a prior distribution $\text{Beta}(1,1)$, corresponding to a uniform prior over $[0,1]$. Given that the label $Y\in\{+1,-1\}$, we model the conditional distribution $Y\mid S=s$ as a Bernoulli random variable. To estimate $p_{s}$, we draw samples $s^{\prime}\sim S$ by first sampling $x\sim{\mathcal{X}}$ and then computing the corresponding anomaly score $s^{\prime}$. We record a success ($b=1$) if $s^{\prime}\leq s$, and a failure ($b=0$) otherwise. Repeating this procedure $n$ times yields $t$ successes and ntn-t failures. Then, according to Theorem 2 in Perini et al., (2021), the posterior distribution of $P_{s}$ given the observed binary outcomes $b_{1},\dots,b_{n}$ is $\text{Beta}(1+t,1+n-t)$. We estimate $p_{s}$ using the posterior mean of $P_{s}$ as
@@ 5 Experiments
We evaluate the effectiveness of EPHAD for unsupervised AD across a range of datasets, including visual AD datasets (Section 5.1), tabular AD datasets (Section 5.2), and an industrial AD use case (Appendix C.2). To systematically investigate the impact of contamination at different levels in a rigorous and reproducible way, we introduce controlled contamination into the data, adhering to the experimental design employed in several prior studies (Jiang et al.,, 2022; Wang et al.,, 2025; Zhou and Wu,, 2024). The evidence functions employed in the experiments are computed in an unsupervised manner without utilising ground-truth labels in the test set $\mathcal{D_{\mathrm{test}}}$, mitigating the risk of overfitting. Unless stated otherwise, we use a contamination factor of $\epsilon=0.1$ and a parameter $\beta=0.5$. An ablation study on different values of $\epsilon$ and $\beta$ is presented in Section 5.3. For image and tabular datasets, we evaluate performance using the AUROC. Following prior work (Roth et al.,, 2022; Gudovskiy et al.,, 2022), AUROC is averaged across all categories for each dataset.
@@ 5.1 Experiments on visual AD datasets
Benchmark datasets. We assess the effectiveness of EPHAD in both sensory and semantic anomaly detection. Sensory AD focuses on detecting physical defects or imperfections, such as a broken capsule or a cut in a carpet, while semantic AD identifies anomalies belonging to a different semantic class—for instance, treating cats as normal and any other animal as anomalous. For sensory AD in industrial contexts, we evaluate performance using four well-established benchmark datasets: MVTecAD (Bergmann et al.,, 2019), MPDD (Jezek et al.,, 2021), ViSA (Zou et al.,, 2022), and RealIAD (Wang et al.,, 2024). For semantic AD, we utilise four commonly used datasets, including CIFAR-$10$, Fashion-MNIST, MNIST, and SVHN. Following the one-vs-rest protocol (Qiu et al.,, 2022), we construct $k$ AD tasks per dataset, where $k$ corresponds to the number of classes. For MVTecAD, ViSA, MPDD and RealIAD, we adopt the “overlap” setting, introducing $\epsilon\%$ contamination into the training set by randomly selecting anomalous samples from the test set while retaining them in the test set Jiang et al., (2022). For the remaining datasets, we follow the “non-overlapping” setting, excluding anomalous samples used for contamination simulation from the test set. Our implementation is based on the public codebase from Jiang et al., (2022). Additional details are provided in Appendix B.1.
@@
Baseline AD methods. We evaluate the performance of several state-of-the-art unsupervised anomaly detection methods, including PatchCore (Roth et al.,, 2022), PaDim (Defard et al.,, 2021), CFLOW (Gudovskiy et al.,, 2022), FastFLOW (Yu et al.,, 2021), DRÆM (Zavrtanik et al.,, 2021), Reverse Distillation (RD) (Deng and Li,, 2022), and ULSAD (Patra and Ben Taieb,, 2024), both with and without the integration of EPHAD. Implementations for all methods, except ULSAD, are based on the Anomalib library (Akcay et al.,, 2022), while ULSAD is implemented using its official public code. Since, to the best of our knowledge, no existing AD method with contaminated data offers post-hoc adaptation in the same manner as EPHAD, our primary objective is to demonstrate the effectiveness of EPHAD by comparing its relative performance against the AD model and the evidence function alone. We also provide comparative analyses with three existing frameworks “Refine” (Yoon et al.,, 2022), Latent Outlier Exposure (LOE) (Qiu et al.,, 2022), and SoftPatch (Jiang et al.,, 2022) in Appendix C.3.
@@
Evidence function. For the experiments, we employ Contrastive Language-Image Pre-training (CLIP) (Radford et al.,, 2021) as the evidence function for image-based datasets, following the anomaly detection approach as in WinCLIP (Jeong et al.,, 2023). We use CLIP as the evidence function $T(x)$ in EPHAD. We start by defining two lists of textual prompt templates, ${\mathcal{T}}_{N}=\{n_{1},\cdots,n_{k}\}$ and ${\mathcal{T}}_{A}=\{a_{1},\cdots,a_{k}\}$, corresponding to normal and anomalous classes, respectively. These templates are dataset-dependent, reflecting subjectivity (e.g., “missing wire" as anomalous for cables). For each label, compute the mean of text embeddings $t_{N}$ and $t_{A}$. Finally, given an input image $x$, the evidence $T(x)$ at test-time is computed as:
$T(x):=\frac{\exp\left(\langle e_{i}(x),t_{A}\rangle/\gamma\right)}{\exp\left(\langle e_{i}(x),t_{N}\rangle/\gamma\right)+\exp\left(\langle e_{i}(x),t_{A}\rangle/\gamma\right)}.$
@@
Results. In our experiments, as we adopt CLIP in the same manner as WinCLIP (Jeong et al.,, 2023), the baseline CLIP results reported here directly correspond to the standalone performance of WinCLIP. In Table 1, we observe that while zero-shot AD using CLIP performs well on real-world image datasets such as CIFAR10 and FMNIST, its effectiveness declines on domain-specific datasets like MVTec, MPDD, and ViSA, where existing AD methods, such as ULSAD, achieve superior performance. However, when these AD methods are used within the EPHAD framework with CLIP as an evidence function in a post-hoc manner, their performance improves in most cases. Notably, even when CLIP-based AD alone does not achieve the best results, as seen in SVHN, incorporating it within EPHAD still leads to significant improvements. For instance, CFLOW, PaDiM, and RD exhibit enhanced performance after using EPHAD, surpassing both CLIP and the standalone AD methods. This highlights the effectiveness of EPHAD in refining anomaly scores for better AD performance. In some cases, such as ULSAD on SVHN, we observe a decline in performance when integrating EPHAD compared to the standalone AD method. This typically occurs when the AD method substantially outperforms the evidence function. In such scenarios, overly relying on the evidence can diminish overall performance. To mitigate this effect, careful tuning of $\beta$ enables the framework to adapt effectively to different datasets, AD methods, and evidence functions. A detailed analysis of the impact of varying $\beta$ values is presented in Section 5.3.
@@ 5.2 Experiments on tabular AD datasets
Benchmark datasets. We evaluate our proposed framework on $26$ classical benchmark datasets from ADBench (Han et al.,, 2022). The classical datasets include datasets from different domains such as healthcare (e.g., annthyroid, breastw), astronautics (e.g. satellite), and finance (fraud). Following Qiu et al., (2022), we preprocess, split the dataset into the train and test sets and simulate contamination using synthetic anomalies created by adding zero-mean Gaussian noise with a large variance to the anomalous sample from the test set.
Baseline AD methods. We compare EPHAD against IFOREST (Liu et al.,, 2012), LOF (Breunig et al.,, 2000), DeepSVDD (Ruff et al.,, 2018), ECOD (Li et al., 2023b, ) and COPOD (Li et al.,, 2020) using ADBench (Han et al.,, 2022).
Evidence function. We use the output of Local Outlier Factor (LOF) (Breunig et al.,, 2000) and Isolation Forest (IForest) (Liu et al.,, 2012). Additional details provided in the Appendix B.3.2.
@@ B.1 Benchmark datasets
For sensory AD in industrial settings, we use three widely recognised benchmark datasets. MVTecAD (Bergmann et al.,, 2019) comprises images from $15$ categories ($10$ objects and $5$ textures) with $3629$ normal training images and $1258$ anomalous and $467$ normal test images, each containing pixel-level annotations of defects. MPDD (Jezek et al.,, 2021) targets metal part defects under varying conditions, offering $888$ training images and test datasets consisting of $176$ normal and $282$ anomalous images across $6$ metal part categories. ViSA (Zou et al.,, 2022) provides $10821$ high-resolution images ($9621$ normal and $1200$ anomalous) spanning $12$ categories, capturing a range of anomalies such as scratches, cracks, missing parts, and misplacements. Each defect type is represented by 15152020 images, and some images feature multiple defects. RealIAD (Wang et al.,, 2024) is a large-scale industrial AD dataset comprising $\sim 150k$ images across $30$ categories and having various types of defects such as scratches, dirt and missing parts. For experiments with RealIAD, we use the training split with $10\%$ contamination and the test split provided by the authors. For the semantic datasets, using the one-vs-rest protocol, we create $k$ AD tasks for each dataset, where $k$ is the number of classes. In each task, one class is designated as normal, while the remaining classes are treated as anomalous. Across both sensory and semantic AD, the training datasets consist of a mixture of normal samples and a fraction $\epsilon$ of anomalous samples, reflecting realistic contamination scenarios.
@@ B.2 Details of the experiment using synthetic Data
The synthetic dataset is generated using a 22D Gaussian mixture model with three components. Normal samples are drawn from $f^{+}_{X}(x):={\mathcal{N}}([1,1]^{T},0.07\mathbf{I}_{2})$, while anomalous samples are sampled from $f^{-}_{X}(x):={\mathcal{N}}([-0.25,2.5]^{T},0.03\mathbf{I}_{2})+{\mathcal{N}}([-1,0.5]^{T},0.03\mathbf{I}_{2})$. For the experiments, we use DeepSVDD with a one-layer radial basis function (RBF) network. The hidden layer comprises three neurons, with their centres fixed at the mean of each Gaussian component, while the scales are optimized during training. The RBF network outputs a 11D scalar obtained as a linear combination of the outputs from the hidden layer. The centre is initialized randomly and made trainable, with an added bias term in the final layer. Although these modifications are not recommended by Ruff et al., (2018) to avoid collapse to a trivial solution, Qiu et al., (2022) observed that these changes enhance model flexibility and convergence. Following this, we train DeepSVDD using the Adam optimizer with a learning rate of $0.01$, $200$ epochs, and a mini-batch size of $25$.

### v2

@@ 4.1 Extension to score-based anomaly detection
Following energy-based model (EBM) formulations (LeCun et al., 2006), we can define the associated contaminated PDF as
@@ 4.2 An illustrative example
To illustrate the effect of EPHAD, we use a toy dataset inspired by Qiu et al. (2022). The dataset is generated using a two-dimensional mixture model comprising three Gaussian components: $c_{1}:={\mathcal{N}}(\mu_{1},\Sigma_{1}),c_{2}:={\mathcal{N}}(\mu_{2},\Sigma_{2}),c_{3}:={\mathcal{N}}(\mu_{3},\Sigma_{3})$. Here, each component follows a Gaussian distribution ${\mathcal{N}}(\mu,\Sigma)$ with mean $\mu$ and covariance $\Sigma$. Normal samples are drawn from $f^{+}_{X}=c_{1}$, with $\mu_{1}=[1,1]^{T}$ and $\Sigma_{1}=0.07~\mathbf{I}_{2}$. Anomalous samples are drawn from a mixture distribution $f^{-}_{X}:=0.5c_{2}+0.5c_{3}$ where $\mu_{2}=[-0.25,2.5]^{T},\mu_{3}=[-1,0.5]^{T}$ and $\Sigma_{2}=\Sigma_{3}=0.03~\mathbf{I}_{2}$. The extended implementation details is provided in Appendix B.2. Using this setting, we create a contaminated dataset consisting $100$ data points. We compare the baseline DeepSVDD (Ruff et al., 2018) across three configurations as illustrated in Figure 1: (i) “Blind”, (ii) “Refine”, and (iii) with EPHAD. We refer to the baseline model that treats all samples as normal as “Blind”, while “Refine” denotes a model that iteratively filters out suspected anomalies during training. As an evidence function in EPHAD, LOF (Breunig et al., 2000) is computed on test samples at test time. The results in Figure 1 demonstrate that the “Blind” configuration mistakenly considers all anomalies as normal. The “Refine” configuration improves performance by filtering out a subset of anomalies. Finally, EPHAD establishes a clearer boundary around normal samples.
@@ 4.3 Determining the temperature parameter $\beta$
As previously discussed, EPHAD has only a single hyperparameter, $\beta$, which controls the trade-off between reliance on the original AD model and the evidence function $T(x)$. A straightforward approach to selecting $\beta$ would involve evaluating the AD performance of the prior and $T(x)$ individually on a validation set and choosing $\beta$ accordingly. However, this strategy introduces additional computational overhead at test time and requires access to a labelled validation set of sufficient size to ensure reliable performance estimation – conditions often impractical in real-world deployments. To address this limitation, we propose an adaptive extension of our approach, termed EPHAD-Ada, that determines the optimal $\beta$ in an unsupervised manner using only test data at test time. This adaptation is inspired by the principle of Entropy Minimisation (EM) (Press et al., 2024), a widely-used technique in test-time adaptation (Xiao and Snoek, 2024). Motivated by the observation from Wang et al. (2021) that models tend to be more accurate when predictions are made with high confidence, we apply it to compute the hyperparameter $\beta$. Specifically, the computation of $\beta$ depends on the entropy of the inlier probabilities derived from the scores of both the original model and the evidence function.
@@
where $p_{s}:={\mathbb{P}}(S\leq s)$. Since $p_{s}$ is unknown in practice, we follow the approach of Perini et al. (2021) and treat it as a random variable $P_{s}$ with a prior distribution $\text{Beta}(1,1)$, corresponding to a uniform prior over $[0,1]$. Given that the label $Y\in\{+1,-1\}$, we model the conditional distribution $Y\mid S=s$ as a Bernoulli random variable. To estimate $p_{s}$, we draw samples $s^{\prime}\sim S$ by first sampling $x\sim{\mathcal{X}}$ and then computing the corresponding anomaly score $s^{\prime}$. We record a success ($b=1$) if $s^{\prime}\leq s$, and a failure ($b=0$) otherwise. Repeating this procedure $n$ times yields $t$ successes and ntn-t failures. Then, according to Theorem 2 in Perini et al. (2021), the posterior distribution of $P_{s}$ given the observed binary outcomes $b_{1},\dots,b_{n}$ is $\text{Beta}(1+t,1+n-t)$. We estimate $p_{s}$ using the posterior mean of $P_{s}$ as
@@ 5 Experiments
We evaluate the effectiveness of EPHAD for unsupervised AD across a range of datasets, including visual AD datasets (Section 5.1), tabular AD datasets (Section 5.2), and an industrial AD use case (Appendix C.2). To systematically investigate the impact of contamination at different levels in a rigorous and reproducible way, we introduce controlled contamination into the data, adhering to the experimental design employed in several prior studies (Jiang et al., 2022; Wang et al., 2025; Zhou and Wu, 2024). The evidence functions employed in the experiments are computed in an unsupervised manner without utilising ground-truth labels in the test set $\mathcal{D_{\mathrm{test}}}$, mitigating the risk of overfitting. Unless stated otherwise, we use a contamination factor of $\epsilon=0.1$ and a parameter $\beta=0.5$. An ablation study on different values of $\epsilon$ and $\beta$ is presented in Section 5.3. For image and tabular datasets, we evaluate performance using the AUROC. Following prior work (Roth et al., 2022; Gudovskiy et al., 2022), AUROC is averaged across all categories for each dataset.
@@ 5.1 Experiments on visual AD datasets
Benchmark datasets. We assess the effectiveness of EPHAD in both sensory and semantic anomaly detection. Sensory AD focuses on detecting physical defects or imperfections, such as a broken capsule or a cut in a carpet, while semantic AD identifies anomalies belonging to a different semantic class—for instance, treating cats as normal and any other animal as anomalous. For sensory AD in industrial contexts, we evaluate performance using four well-established benchmark datasets: MVTecAD (Bergmann et al., 2019), MPDD (Jezek et al., 2021), ViSA (Zou et al., 2022), and RealIAD (Wang et al., 2024). For semantic AD, we utilise four commonly used datasets, including CIFAR-$10$, Fashion-MNIST, MNIST, and SVHN. Following the one-vs-rest protocol (Qiu et al., 2022), we construct $k$ AD tasks per dataset, where $k$ corresponds to the number of classes. For MVTecAD, ViSA, MPDD and RealIAD, we adopt the “overlap” setting, introducing $\epsilon\%$ contamination into the training set by randomly selecting anomalous samples from the test set while retaining them in the test set Jiang et al. (2022). For the remaining datasets, we follow the “non-overlapping” setting, excluding anomalous samples used for contamination simulation from the test set. Our implementation is based on the public codebase from Jiang et al. (2022). Additional details are provided in Appendix B.1.
@@
Baseline AD methods. We evaluate the performance of several state-of-the-art unsupervised anomaly detection methods, including PatchCore (Roth et al., 2022), PaDim (Defard et al., 2021), CFLOW (Gudovskiy et al., 2022), FastFLOW (Yu et al., 2021), DRÆM (Zavrtanik et al., 2021), Reverse Distillation (RD) (Deng and Li, 2022), and ULSAD (Patra and Ben Taieb, 2024), both with and without the integration of EPHAD. Implementations for all methods, except ULSAD, are based on the Anomalib library (Akcay et al., 2022), while ULSAD is implemented using its official public code. Since, to the best of our knowledge, no existing AD method with contaminated data offers post-hoc adaptation in the same manner as EPHAD, our primary objective is to demonstrate the effectiveness of EPHAD by comparing its relative performance against the AD model and the evidence function alone. We also provide comparative analyses with three existing frameworks “Refine” (Yoon et al., 2022), Latent Outlier Exposure (LOE) (Qiu et al., 2022), and SoftPatch (Jiang et al., 2022) in Appendix C.3.
@@
Evidence function. For the experiments, we employ Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) as the evidence function for image-based datasets, following the anomaly detection approach as in WinCLIP (Jeong et al., 2023). We use CLIP as the evidence function $T(x)$ in EPHAD. We start by defining two lists of textual prompt templates, ${\mathcal{T}}_{N}=\{n_{1},\cdots,n_{k}\}$ and ${\mathcal{T}}_{A}=\{a_{1},\cdots,a_{k}\}$, corresponding to normal and anomalous classes, respectively. These templates are dataset-dependent, reflecting subjectivity (e.g., “missing wire" as anomalous for cables). For each label, compute the mean of text embeddings $t_{N}$ and $t_{A}$. Finally, given an input image $x$, the evidence $T(x)$ at test-time is computed as:
$T(x):=\frac{\exp\left(\langle e_{i}(x),t_{N}\rangle/\gamma\right)}{\exp\left(\langle e_{i}(x),t_{N}\rangle/\gamma\right)+\exp\left(\langle e_{i}(x),t_{A}\rangle/\gamma\right)}.$
@@
Results. In our experiments, as we adopt CLIP in the same manner as WinCLIP (Jeong et al., 2023), the baseline CLIP results reported here directly correspond to the standalone performance of WinCLIP. In Table 1, we observe that while zero-shot AD using CLIP performs well on real-world image datasets such as CIFAR10 and FMNIST, its effectiveness declines on domain-specific datasets like MVTec, MPDD, and ViSA, where existing AD methods, such as ULSAD, achieve superior performance. However, when these AD methods are used within the EPHAD framework with CLIP as an evidence function in a post-hoc manner, their performance improves in most cases. Notably, even when CLIP-based AD alone does not achieve the best results, as seen in SVHN, incorporating it within EPHAD still leads to significant improvements. For instance, CFLOW, PaDiM, and RD exhibit enhanced performance after using EPHAD, surpassing both CLIP and the standalone AD methods. This highlights the effectiveness of EPHAD in refining anomaly scores for better AD performance. In some cases, such as ULSAD on SVHN, we observe a decline in performance when integrating EPHAD compared to the standalone AD method. This typically occurs when the AD method substantially outperforms the evidence function. In such scenarios, overly relying on the evidence can diminish overall performance. To mitigate this effect, careful tuning of $\beta$ enables the framework to adapt effectively to different datasets, AD methods, and evidence functions. A detailed analysis of the impact of varying $\beta$ values is presented in Section 5.3.
@@ 5.2 Experiments on tabular AD datasets
Benchmark datasets. We evaluate our proposed framework on $26$ classical benchmark datasets from ADBench (Han et al., 2022). The classical datasets include datasets from different domains such as healthcare (e.g., annthyroid, breastw), astronautics (e.g. satellite), and finance (fraud). Following Qiu et al. (2022), we preprocess, split the dataset into the train and test sets and simulate contamination using synthetic anomalies created by adding zero-mean Gaussian noise with a large variance to the anomalous sample from the test set.
Baseline AD methods. We compare EPHAD against IFOREST (Liu et al., 2012), LOF (Breunig et al., 2000), DeepSVDD (Ruff et al., 2018), ECOD (Li et al., 2023b) and COPOD (Li et al., 2020) using ADBench (Han et al., 2022).
Evidence function. We use the output of Local Outlier Factor (LOF) (Breunig et al., 2000) and Isolation Forest (IForest) (Liu et al., 2012). Additional details provided in the Appendix B.3.2.
@@ B.1 Benchmark datasets
For sensory AD in industrial settings, we use three widely recognised benchmark datasets. MVTecAD (Bergmann et al., 2019) comprises images from $15$ categories ($10$ objects and $5$ textures) with $3629$ normal training images and $1258$ anomalous and $467$ normal test images, each containing pixel-level annotations of defects. MPDD (Jezek et al., 2021) targets metal part defects under varying conditions, offering $888$ training images and test datasets consisting of $176$ normal and $282$ anomalous images across $6$ metal part categories. ViSA (Zou et al., 2022) provides $10821$ high-resolution images ($9621$ normal and $1200$ anomalous) spanning $12$ categories, capturing a range of anomalies such as scratches, cracks, missing parts, and misplacements. Each defect type is represented by 15152020 images, and some images feature multiple defects. RealIAD (Wang et al., 2024) is a large-scale industrial AD dataset comprising $\sim 150k$ images across $30$ categories and having various types of defects such as scratches, dirt and missing parts. For experiments with RealIAD, we use the training split with $10\%$ contamination and the test split provided by the authors. For the semantic datasets, using the one-vs-rest protocol, we create $k$ AD tasks for each dataset, where $k$ is the number of classes. In each task, one class is designated as normal, while the remaining classes are treated as anomalous. Across both sensory and semantic AD, the training datasets consist of a mixture of normal samples and a fraction $\epsilon$ of anomalous samples, reflecting realistic contamination scenarios.
@@ B.2 Details of the experiment using synthetic Data
The synthetic dataset is generated using a 22D Gaussian mixture model with three components. Normal samples are drawn from $f^{+}_{X}(x):={\mathcal{N}}([1,1]^{T},0.07\mathbf{I}_{2})$, while anomalous samples are sampled from $f^{-}_{X}(x):={\mathcal{N}}([-0.25,2.5]^{T},0.03\mathbf{I}_{2})+{\mathcal{N}}([-1,0.5]^{T},0.03\mathbf{I}_{2})$. For the experiments, we use DeepSVDD with a one-layer radial basis function (RBF) network. The hidden layer comprises three neurons, with their centres fixed at the mean of each Gaussian component, while the scales are optimized during training. The RBF network outputs a 11D scalar obtained as a linear combination of the outputs from the hidden layer. The centre is initialized randomly and made trainable, with an added bias term in the final layer. Although these modifications are not recommended by Ruff et al. (2018) to avoid collapse to a trivial solution, Qiu et al. (2022) observed that these changes enhance model flexibility and convergence. Following this, we train DeepSVDD using the Adam optimizer with a learning rate of $0.01$, $200$ epochs, and a mini-batch size of $25$.